# Private Benchmarking to Prevent Contamination and Improve Comparative Evaluation of LLMs

### v1

@@ 3.2.2 EzPC technology to evaluate models on benchmarks (P2P)
EzPC Chandran et al. (2019); Rathee et al. (2020); Jawalkar et al. (2024); Gupta et al. (2023); team is a cryptographically secure system built on the basis of the seminal cryptographic technology of secure multi-party computation Goldreich et al. (1987); Yao (1986). It allows multiple entities to bring in private data and enable them to perform computations on the data without trusting any entity or each other in the computation. This is done through an interactive protocol in which parties perform computations as well as interact and exchange (seemingly random looking) messages with the other party. The security guarantees provided are cryptographic (i.e. unless cryptography is broken, e.g. AES-128 is broken, no entity can learn any information other than the results of the computation, which in this case would be the evaluation results).
@@
EzPC technology enables secure computation of LLMs and provides support for execution on CPUs as well as GPUs. It has also been developed and tested in the context of model evaluation in the healthcare domain (see Soin et al. (2021)). Secure multi-party computation based solutions such as EzPC assume that the model architecture is known to all participants; while current efficient systems (such as the EzPC system) typically assume that the participants execute the protocol code given to them faithfully (called semi-honest behavior in cryptography), solutions also exist that avoid such an assumptions Keller (2020, 2019). The solution architecture for this case is illustrated in Figure 5.
@@ 4 Auditing Datasets
One of the crucial aspects of benchmark creation is ensuring that the benchmark is of high quality and can reliably test model capabilities of interest. To create a representative evaluation dataset, several factors must be taken into consideration, including the source of the data, data selection and filtering practices, as well as annotator demographics and language background Doğruöz et al. (2023). Neglecting these factors can result in low-quality datasets that provide misleading results. Therefore, it is crucial for us to be able to audit benchmarks to ensure that they are of sufficiently high quality. In the case of open benchmarks, this is relatively easy to do.
@@
Honest Benchmark Owner. In this, we assume that the benchmark owner is an honest entity. That is, the benchmark owner curates a benchmark to be presented to the auditor. While this benchmark may not be of a high quality, the owner will not act maliciously to fool an auditor into believing otherwise. In this case, the following method can be used to audit a benchmark. As before, suppose the benchmark has data points with labels . Now, the auditor samples indices at random from and sends these indices to the benchmark owner. The benchmark owner provides the auditor with . We also assume that the test can be performed individually on each benchmark data point to determine its quality (i.e., returns or depending on whether the benchmark is of good quality or not). Now, the auditor performs the test on each of these points and declares the benchmark to be of a good quality if for all . One can then easily show a theorem of the following form: “If the auditor declares the benchmark to be of a good quality, then at least of the dataset is of good quality, except with probability ". For example, if we wanted a guarantee that of the dataset is of a good quality, then by checking points at random, we can ascertain that this is indeed the case, except with probability . Of course, this solution will reveal 100 data points to an auditor and is only applicable when this is acceptable to the benchmark owner.
@@
Somewhat Honest Benchmark Owner. One can also design a solution when we do not wish to assume that the benchmark owner is fully honest as in the above solution. Suppose the benchmark owner curates a benchmark to be presented to the auditor, but the owner may also maliciously try to fool the auditor into believing that this benchmark if of high quality. One can provide a solution in this case as follows. As above, we assume that the benchmark has data points with labels and as above, we assume that the test can be performed individually on each benchmark data point to determine its quality (i.e., returns or depending on whether the benchmark is of good quality or not). In this solution, the benchmark owner will provide a cryptographic commitment Goldreich (2000); Merkle (1987) to each of the data points (a cryptographic commitment to a value can be thought of as providing a locked box that contains the value . The commitment (locked box) hides the value and can be opened later (with an opening or a key). The value inside a locked box cannot be changed (called binding) once sent to the recipient). Once the benchmark owner provides a cryptographic commitment to every , the auditor then does as in the previous solution. That is, the auditor samples indices at random from and sends these indices to the benchmark owner. The benchmark owner will now provide the openings (keys) to each of the commitments (i.e., for ). The auditor will open these commitments, learn the data points and perform the test on these points. The guarantees provided are as before - i.e., one can easily show “If the auditor declares the benchmark to be of a good quality, then at least of the dataset is of good quality, except with probability ".
@@ 5 Conclusion
We present a unique inter-disciplinary solution to this problem - private benchmarking, in which the benchmark datasets are kept private from the model, using advances in secure computation and cryptography. We present solutions to several trust models and describe their pros and cons. The proposed solution has the potential to create cross-industry evaluation platforms like the SPEC benchmarks in computer architecture 555https://www.spec.org/30th/ since corporations can now share their benchmarks for evaluation without compromising on their competitive advantage. Another possibility is that of startups offering secure benchmarking as a service using the provably secure techniques described in this paper (such as in Figure 5) for keeping the model weights and the benchmark data private. This is similar to the creation of companies at the start of the RSA security certification process in the early 90s and the emergence of large companies like Verisign. Another example is that of the Truste seal 666https://www.trustsignals.com/blog/history-of-truste used to signal privacy practices of e-commerce websites in the early 2000s.

### v2

@@ 3.2.2 EzPC technology to evaluate models on benchmarks (P2P)
EzPCΒ [ezpc, cryptflow2, orca, sigma, ezpc-code] is a cryptographically secure system built on the basis of the seminal cryptographic technology of secure multi-party computationΒ [gmw, yaogc]. It allows multiple entities to bring in private data and enable them to perform computations on the data without trusting any entity or each other in the computation. This is done through an interactive protocol in which parties perform computations as well as interact and exchange (seemingly random looking) messages with the other party. The security guarantees provided are cryptographic (i.e. unless cryptography is broken, e.g. AES-128 is broken, no entity can learn any information other than the results of the computation, which in this case would be the evaluation results).
@@
EzPC technology enables secure computation of LLMs and provides support for execution on CPUs as well as GPUs. It has also been developed and tested in the context of model evaluation in the healthcare domain (seeΒ [secureeval]). Secure multi-party computation based solutions such as EzPC assume that the model architecture is known to all participants. While the most efficient systems today (such as the EzPC system) typically assume that the participants execute the protocol code given to them faithfully (called semi-honest behavior in cryptography), solutions also exist that avoid such an assumptions Β [mpspdz, mpspdz-code]. The solution architecture for this case is illustrated in FigureΒ 5. While solutions based on cryptographic secure multiparty computation suffered from astronomical overheads in the past, the latency overheads of EzPC are tractableΒ [sigma].
4 System and Evaluation
4.1 Implementation
We have implemented an open-source platform, called , to help model owners and dataset owners to collaborate and create public leaderboards. For evaluation, the platform supports all five benchmarking options: trusted model owner (FigureΒ 2), trusted dataset owner (FigureΒ 2), trusted third party (FigureΒ 4), trusted execution environment (FigureΒ 4), and cryptographic secure multiparty computation with EzPC (FigureΒ 5). The platform is implemented in 5k lines of python using Django as the backend.
A model owner can raise a request to evaluate the available private datasets using any of the five benchmarking methods. Conversely, this mandates that both the model owner and dataset owner agree to run the evaluation. For each type of evaluation submission, the platform will generate a unique ephemeral certificate, signed by the platformβs Certificate Authority, for that submission only. This certificate is then used to ensure encrypted communication using TLSv1.3 and authenticate parties to each other in TEE, TTP (third party), and EzPC (P2P). After the evaluation is complete, the results are pushed to the platformβs public leaderboard.
The platform is publicly available at github.com/microsoft/private-benchmarking under MIT license.
4.2 Evaluation
| Model + Dataset | Type | Time per sample (s) | #samples | Total Communication (GB) |
| Bert-BaseΒ [devlin2018bert] | TTP | 0.00137 | 0.40 | |
| on | TEE | 0.00161 | 1,821 | 0.40 |
| SST2Β [socher-etal-2013-recursive] | EzPC | 1.322 | 1,442.13 | |
| AiravataΒ [gala2024airavata] | TTP | 0.00832 | 25.60 | |
| on | TEE | 0.00845 | 5,010 | 25.60 |
| IndicxnliΒ [aggarwal-etal-2022-indicxnli] | EzPC | 12.157 | 75,711.03 | |
| Llama2-7BΒ [touvron2023llama] | TTP | 0.05298 | 25.53 | |
| on | TEE | 0.05354 | 5153 | 25.53 |
| LambadaΒ [lambada_dataset] | EzPC | 12.269 | 1,98,996.58 | |
| VGG16Β [simonyan2015deep] | TTP | 0.00055 | 0.69 | |
| on | TEE | 0.00058 | 10,000 | 0.69 |
| ImageNet-1kΒ [russakovsky2015imagenet] | EzPC | 0.329 | 3,166.97 |
We evaluate the time and communication of various private benchmarking options on publicly available models and datasets in TableΒ 1. We find that compared to the existing baseline solutions (SectionΒ 3.1), the overheads of confidential computing (using confidential Hopper GPUs) are negligible and that of EzPC (using SigmaΒ [sigma]) are tractable.
Evaluation setup.
We evaluate on Azure virtual machines (VMs) where each VM has an Nvidiaβs Hopper H100 GPU, 40 CPU cores, and 320 GB RAM. To benchmark using a trusted third party (TTP), both the model and the dataset are loaded into a single VM which runs the evaluation on the GPU. For evaluation with trusted execution environment, the GPU is put in the confidential mode. For evaluation with EzPC, we use two such VMs that are connected via a computer network with 24Gbps bandwidth.
Performance.
The baseline TTP time for the language models is obtained by running ten samples with 100 tokens each and taking the mean of the runtimes. For the vision model VGG16, we run it on ten images and take the mean. We follow the same process for TEEs and the running time increases slightly. The running time of cryptographic computation with EzPC is higher but tractable. The variations in runtime of the different mechanisms are negligible.
Apart from time, another metric of interest is communication. The TTP and TEE options must communicate the model and the dataset to the evaluating VM. Subsequently, no communication is required during the evaluation process. With EzPC, the VM belonging to the model owner and the VM belonging to the dataset owner must communicate encrypted data for each inference. While the communication of EzPC is much higher, cryptographic solutions are becoming more communication frugal with each passing yearΒ [ezpc, cryptflow2, orca, sigma]. Note that the time per sample reported for EzPC in TableΒ 1 includes the time taken in communicating the encryptions as well.
@@ 5 Auditing Datasets
One of the crucial aspects of benchmark creation is ensuring that the benchmark is of high quality and can reliably test model capabilities of interest. To create a representative evaluation dataset, several factors must be taken into consideration, including the source of the data, data selection and filtering practices, as well as annotator demographics and language background [dogruoz-etal-2023-representativeness]. Neglecting these factors can result in low-quality datasets that provide misleading results. Therefore, it is crucial for us to be able to audit benchmarks to ensure that they are of sufficiently high quality. In the case of open benchmarks, this is relatively easy to do. Similarly, owners of large datasets can randomly partition their datasets into training sets which are released publicly and keep small test sets privateΒ [gaia].
@@
Honest Benchmark Owner. In this, we assume that the benchmark owner is an honest entity. That is, the benchmark owner curates a benchmark to be presented to the auditor. While this benchmark may not be of a high quality, the owner will not act maliciously to fool an auditor into believing otherwise. In this case, the following method can be used to audit a benchmark. As before, suppose the benchmark has data points with labels . Now, the auditor samples indices at random from and sends these indices to the benchmark owner. The benchmark owner provides the auditor with . We also assume that the test can be performed individually on each benchmark data point to determine its quality (i.e., returns or depending on whether the benchmark is of good quality or not). Now, the auditor performs the test on each of these points and declares the benchmark to be of a good quality if for all . One can then easily show a theorem of the following form: βIf the auditor declares the benchmark to be of a good quality, then at least of the dataset is of good quality, except with probability ". For example, if we wanted a guarantee that of the dataset is of a good quality, then by checking points at random, we can ascertain that this is indeed the case, except with probability . Of course, this solution will reveal 100 data points to an auditor and is only applicable when this is acceptable to the benchmark owner.
@@
Somewhat Honest Benchmark Owner. One can also design a solution when we do not wish to assume that the benchmark owner is fully honest as in the above solution. Suppose the benchmark owner curates a benchmark to be presented to the auditor, but the owner may also maliciously try to fool the auditor into believing that this benchmark if of high quality. One can provide a solution in this case as follows. As above, we assume that the benchmark has data points with labels and as above, we assume that the test can be performed individually on each benchmark data point to determine its quality (i.e., returns or depending on whether the benchmark is of good quality or not). In this solution, the benchmark owner will provide a cryptographic commitmentΒ [goldreichbook, merklecommitment] to each of the data points (a cryptographic commitment to a value can be thought of as providing a locked box that contains the value . The commitment (locked box) hides the value and can be opened later (with an opening or a key). The value inside a locked box cannot be changed (called binding) once sent to the recipient). Once the benchmark owner provides a cryptographic commitment to every , the auditor then does as in the previous solution. That is, the auditor samples indices at random from and sends these indices to the benchmark owner. The benchmark owner will now provide the openings (keys) to each of the commitments (i.e., for ). The auditor will open these commitments, learn the data points and perform the test on these points. The guarantees provided are as before - i.e., one can easily show βIf the auditor declares the benchmark to be of a good quality, then at least of the dataset is of good quality, except with probability ".
@@ 6 Conclusion
We present a unique inter-disciplinary solution to this problem - private benchmarking, in which the benchmark datasets are kept private from the model, using advances in secure computation and cryptography. We present solutions to several trust models, describe their pros and cons, build an end-to-end platform for such private benchmarking and compare the various solutions on performance aspects. Another possibility is that of startups offering private benchmarking as a service using the provably secure techniques described in this paper (such as in Figure 5) for keeping the model weights and the benchmark data private. This is similar to the creation of companies at the start of the RSA security certification process in the early 90s and the emergence of large companies like Verisign. Another example is that of the Truste seal 555https://www.trustsignals.com/blog/history-of-truste used to signal privacy practices of e-commerce websites in the early 2000s.
@@ 7 Limitations
Acknowledgements
We thank Kapil Vaswani and Krishnaprasad Hande for providing us access to confidential GPUs. We thank Neha Jawalkar and Kanav Gupta for help with EzPC.